# Cross-Modal Bidirectional Interaction Model for Referring Remote Sensing Image Segmentation

### v1

@@ Abstract
Given a natural language expression and a remote sensing image, the goal of referring remote sensing image segmentation (RRSIS) is to generate a pixel-level mask of the target object identified by the referring expression. In contrast to natural scenarios, expressions in RRSIS often involve complex geospatial relationships, with target objects of interest that vary significantly in scale and lack visual saliency, thereby increasing the difficulty of achieving precise segmentation. To address the aforementioned challenges, a novel RRSIS framework is proposed, termed the cross-modal bidirectional interaction model (CroBIM). Specifically, a context-aware prompt modulation (CAPM) module is designed to integrate spatial positional relationships and task-specific knowledge into the linguistic features, thereby enhancing the ability to capture the target object. Additionally, a language-guided feature aggregation (LGFA) module is introduced to integrate linguistic information into multi-scale visual features, incorporating an attention deficit compensation mechanism to enhance feature aggregation. Finally, a mutual-interaction decoder (MID) is designed to enhance cross-modal feature alignment through cascaded bidirectional cross-attention, thereby enabling precise segmentation mask prediction. To further forster the research of RRSIS, we also construct RISBench, a new large-scale benchmark dataset comprising 52,472 image-language-label triplets. Extensive benchmarking on RISBench and two other prevalent datasets demonstrates the superior performance of the proposed CroBIM over existing state-of-the-art (SOTA) methods. The source code for CroBIM and the RISBench dataset will be publicly available at https://github.com/HIT-SIRS/CroBIM .
Index Terms:
@@
I Introduction
Over the past few years, deep learning has emerged as the cornerstone for a diverse range of remote sensing applications. Early efforts in intelligent interpretation within the remote sensing domain largely centered on the extraction of visual features from imagery to perform various tasks such as semantic segmentation[1], object detection[2], and change detection[3]. Despite the significant progress made, these studies have predominantly concentrated on visual comprehension, frequently neglecting the critical aspect of modeling object relationships and achieving a more profound semantic understanding.
Recently, large language models (LLMs) have achieved unprecedented advancements in language comprehension, driven by their extensive expert knowledge and sophisticated reasoning abilities. This progress has, in turn, catalyzed significant research into vision-language models (VLMs). The integration of natural language with remote sensing imagery has emerged as a prominent research focus, encompassing tasks such as image captioning[4, 5], image-text retrieval[6, 7], text-based remote sensing image generation[8, 9], and visual question answering[10, 11]. Despite these advancements, the task of referring remote sensing image segmentation (RRSIS) remains relatively unexplored.
@@ I Introduction
    We present a novel framework for the RRSIS task, named CroBIM, designed to address the significant challenges posed by the diversity of remote sensing data and the complex geospatial relationships inherent in the corresponding expressions.
@@
    We design the CAPM module to integrate multi-scale visual contextual with linguistic features by introducing learnable prompts, enabling precise recognition of the spatial structure and positioning of target objects. Meanwhile, the LGFA module is proposed to facilitate interaction between visual and linguistic features across multiple scales, capturing cross-scale dependencies and improving feature aggregation through attention mechanisms. Besides, We introduce the MID to achieve precise alignment between vision and language modalities via cascaded bidirectional cross-attention, leading to accurate segmentation mask predictions.
@@
    To foster the research of RRSIS, we meticulously construct the largest benchmark dataset to date, named RISBench. RISBench consists of 52,472 high-quality image-language-label triplets, featuring diverse referring expressions and corresponding masks generated semi-automatically.
@@
The remainder of this paper is structured as follows: Section II reviews related works on RRSIS. Section III details the construction process of our proposed RISBench dataset and provides an analysis of its key characteristics. In Section IV, we describe the proposed methodology in detail. Section V presents a comprehensive set of experiments and in-depth analyses. Finally, Section VII concludes the paper and offers insights into potential future research directions.
II Related Work
II-A Referring Image Segmentation
@@
Hu et al.[26] proposed an innovative approach for referring image segmentation by integrating the convolutional neural network (CNN) and long short-term memory network (LSTM) framework. This approach effectively extracts visual features from images and linguistic features from natural language expressions, enabling precise and accurate image segmentation. A recurrent refinement network (RRN)[27] was proposed to capture multi-scale semantics in image representations. The RRN iteratively optimized the initial mask using a recursive optimization module to achieve a high-quality pixel segmentation mask.
However, the aforementioned methods only focus on a single modality of vector representation, neglecting the modality gap and not fully considering the complex interaction between language expressions and images. To address the aforementioned limitations, attention mechanisms have been introduced in recent works. A cross-modal self-attention (CMSA) module by Ye et al.[28] was proposed to effectively captures long-range dependencies between language and visual features. A cascade-grouped attention network (CGAN)[29] is designed, consisting of cascade-grouped attention (CGA) and instance-level attention loss (ILA). By performing hierarchical reasoning on images and effectively distinguishing different instances, CGAN enhances the correlation between text and images. Besides, Hu et al.[30] introduced a bidirectional relationship inferring network (BRINet) to model cross-modal information dependencies. BRINet utilized a visual-guided language attention module to filter out irrelevant regions and enhance semantic matching between target objects and expressions.
II-B Visual Grounding for Remote Sensing Data
@@
II-C Referring Remote Sensing Image Segmentation
@@
TABLE I: Comparative Analysis of RISBench Dataset and Previous Datasets.
Dataset Num. of Triplets Image Size Spatial Resolution Attributes of Expression
RefSegRS[18] 4,420 512x512 0.13m 3
RRSIS-D[20] 17,402 800x800 0.5m - 30m 7
RISBench 52,472 512x512 0.1m - 30m 8
@@
III Dataset Construction and Analysis
In this section, we will introduce the construction procedure and statistical analysis of our proposed RISBench dataset in Section III-A and Section III-B.
III-A RISBench Dataset Construction
  • Step 2. To ensure the accuracy and reliability of the generated mask annotations, we employ the PA-SAM method [34] to optimize the SAM model. This enhancement process focuses on improving the quality of the segmentation masks by addressing the deficiencies in the initial bounding boxes. Specifically, for each given bounding box, we train a specialized adapter based on box prompts to refine the mask decoder features. This training process involves fine-tuning the SAM model to better handle the nuances and complexities of remote sensing imagery. By optimizing the mask decoder features, we significantly improve SAM’s segmentation performance, resulting in high-quality masks that are better aligned with the visual grounding boxes and accurately represent the segmented objects.
@@ III-A RISBench Dataset Construction
III-B Dataset Statistics
After meticulously filtering out uninformative image-language label triplets, we curated the RISBench dataset, comprising 52,472 high-quality image-language label triplets. This dataset is partitioned into a training set with 26,300 triplets, a validation set with 10,013 triplets and a test set with 16,158 triplets, ensuring robust model development and evaluation. Each image in RISBench is uniformly sized at 512$\times$512 pixels, maintaining consistency across the dataset. The spatial resolution of the images spans from 0.1m to 30m, encompassing a diverse range of scales and details. The semantic labels are categorized into 26 distinct classes, each annotated with 8 attributes, thereby facilitating a comprehensive and nuanced semantic segmentation analysis. As shown in Table. I, compared to the previous RRSIS datasets, our dataset demonstrates significant improvements in both quantity and diversity.
@@ IV Methodology
IV-A Context-Aware Prompt Modulation
@@
IV-B Language-Guided Feature Aggregation
@@
Subsequently, we calculate the attention deficit map $\mathcal{M}\in\mathbb{R}^{H_{4}\times W_{4}}$ between cross-scale attentions and select the top $K$ regions with the highest correlation differences:
@@
IV-C Mutual-Interaction Decoder
@@
Integrating the complementary information between cross-modal features is a fundamental challenge in RRSIS task. Cross-modal features often encompass inconsistent information, and without considering the intermediate interactions and thorough alignment between different modalities, it is impossible to ensure the discriminative power of the learned representations. To address the aforementioned challenges, we have meticulously designed the mutual-interaction decoder (MID), as illustrated in Fig. 8, aiming to achieve more comprehensive cross-modal alignment and precise pixel prediction.
The MID utilizes visual features $\left\{V_{il}\right\}_{i=1}^{4}$ and linguistic features $L_{v}$ as inputs to predict the mask of the referred object. Prior to executing cross-modal alignment, a series of operations are performed to harmonize the dimensions of the visual and linguistic features.
@@
IV-D Training Objective
@@ V Experiments
V-A Dataset and Evaluation Metrics
@@
RISBench includes a total of 52,472 image-language-label triplets. It is divided into two subsets: a training set consisting of 26,300 triplets, a validation set consisting of 10,013 triplets and a test set comprising 16,159 triplets. All images are uniformly formatted to a resolution of 512$\times$512 pixels, with spatial resolutions ranging from 0.1 meters to 30 meters. The dataset’s semantic labels are divided into 26 unique classes, with each class further annotated by 8 attributes.
@@
V-B Experimental Setup
We employ Swin Transformer [35] and ConvNeXt [36] as the visual backbones in our approach. The Swin Transformer backbone is initialized with classification weights from the Swin-Base model pre-trained on ImageNet22K [52], while the ConvNeXt backbone utilizes pre-trained weights from ConvNeXt-Base, obtained through self-supervised learning using the SMLFR algorithm [53]. For the language backbone, we use the base BERT model with 12 layers and a hidden size of 768, as available in the HuggingFace library [54]. The maximum sequence length for text descriptions is set to 20 tokens.
@@ V-B Experimental Setup
TABLE II: Comparison with state-of-the-art methods on the RRSIS-D dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@
TABLE III: Comparison with state-of-the-art methods on the RefSegRS dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@
Figure 9: Visualization of segmentation results for CroBIM and comparison methods on the RRSIS-D dataset test set, with corresponding IoU scores displayed.
Figure 10: Visualization of segmentation results for CroBIM and comparison methods on the RefSegRS dataset test set, with corresponding IoU scores displayed.
V-C Comparison with State-of-the-art Methods
V-C1 RRSIS-D
To evaluate the effectiveness of our proposed method, we conducted experiments on the RRSIS-D dataset. The comparison results are presented in Table. II. We compared CroBIM with several LSTM-based, CLIP-based, and BERT-based methods from classical to state-of-the-art, i.e., RRN[27], CSMA[28], LSCM[41], CMPC[42], BRINet[30], CMPC+[43], BKINet[44], ETRIS[45], CRIS[46], LGCE[18], LAVT[19], RMSIN[20], CrossVLT[47], RIS-DMMI[48], robust-ref-seg[49], SLViT[56], CARIS[50].
V-C2 RefSegRS
We further conduct experiments on the RefSegRS dataset to validate the superiority of our proposed framework, and the performance on both the validation and test sets is reported in Table. III. First, in terms of the Pr@0.5 and Pr@0.6 metrics, CroBIM consistently outperforms other state-of-the-art methods on both the validation and test sets. Notably, CroBIM with ConvNeXt-B as the visual encoder achieves the highest accuracy of 75.89% on the test set for Pr@0.5, significantly surpassing the second-best method, RIS-DMMI, which attained 63.89%. This result demonstrates that CroBIM effectively segments the target regions under lower overlap thresholds. Additionally, CroBIM continues to exhibit a marked advantage in the Pr@0.6 metric, achieving an accuracy of 61.42% on the test set, which is substantially higher than RIS-DMMI’s 44.30%. As the overlap threshold increases, CroBIM maintains robust performance. Specifically, at the more stringent Pr@0.9 threshold, CroBIM attains an accuracy of 6.26% on the test set, further confirming its effectiveness in high-precision segmentation tasks.
TABLE IV: Comparison with state-of-the-art methods on the RISBench dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@ V-C2 RefSegRS
Figure 11: Visualization of segmentation results for CroBIM and comparison methods on the RisBench dataset test set, with corresponding IoU scores displayed
V-C3 RisBench
In our constructed RISBench dataset, the proposed CroBIM model also demonstrates significant performance advantages over competing methods, achieving either the best or second-best results across multiple evaluation metrics, as shown in Table. IV. Notably, CroBIM outperforms the current state-of-the-art models on both the validation and test sets across various threshold levels, including Pr@0.5, Pr@0.6, Pr@0.7, Pr@0.8, and Pr@0.9. These results further validate the effectiveness and robustness of our model. Specifically, the CroBIM model, utilizing the ConvNeXt-B visual encoder and BERT text encoder, achieves the highest test set precision at Pr@0.5, reaching 77.55%, outperforming all other methods. Additionally, at Pr@0.6, Pr@0.7, and Pr@0.8, the model attains precisions of 72.83%, 66.38%, and 55.93%, respectively, significantly surpassing existing methods. Particularly, at the higher threshold of Pr@0.9, CroBIM leads other methods with a precision of 34.07%, demonstrating the model’s strong capability to handle more complex scenarios that demand higher precision.

### v2

@@ Abstract
Referring remote sensing image segmentation (RRSIS) aims to generate pixel-level masks of target objects in remote sensing images based on natural language expressions. This task is inherently challenging due to the complex geospatial relationships and varying object scales in remote sensing scenarios. To address the aforementioned challenges, a novel RRSIS framework is proposed, termed the cross-modal bidirectional interaction model (CroBIM). Specifically, a context-aware prompt modulation (CAPM) module is designed to integrate spatial positional relationships and task-specific knowledge into the linguistic features, thereby enhancing the ability to capture the target object. Additionally, a language-guided feature aggregation (LGFA) module is introduced to incorporate linguistic cues into multi-scale visual features, with an attention deficit compensation mechanism being applied to ensure robust alignment across object scales. Finally, a mutual-interaction decoder (MID) is developed to align vision and language features through iterative bidirectional refinement, enabling precise segmentation even in complex geospatial contexts. Moreover we also construct RISBench, a new large-scale benchmark dataset comprising 52,472 image-language-label triplets. Extensive benchmarking on RISBench and two other prevalent datasets demonstrates the superior performance of the proposed CroBIM over existing state-of-the-art (SOTA) methods. The source code for CroBIM and the RISBench dataset will be publicly available at https://github.com/HIT-SIRS/CroBIM.
keywords:
@@
journal: ISPRS Journal of Photogrammetry and Remote Sensingfootnotetext: * These authors contributed equally to this work.footnotetext: Corresponding author. E-mail address: guyf@hit.edu.cn (Y. Gu).
1 Introduction
Deep learning has become a cornerstone for remote sensing applications[13, 12], enabling tasks such as semantic segmentation [9], object detection[2], and change detection[3]. However, these methods often focus on visual comprehension, neglecting the modeling of object relationships and deeper semantic understanding.
Recent advancements in large language models (LLMs) have spurred research into vision-language models (VLMs), enabling tasks such as image captioning[66, 62], image-text retrieval[1, 57], text-based remote sensing image generation[49, 61], and visual question answering[56, 63]. However, the task of referring remote sensing image segmentation (RRSIS) remains relatively unexplored.
@@ 1 Introduction
    We propose CroBIM, a novel framework specifically designed to address the unique challenges of referring remote sensing image segmentation (RRSIS), including the diversity of remote sensing data and the complexity of geospatial relationships described in natural language.
@@
    We design three task-specific modules to address the core challenges of RRSIS: the context-aware prompt modulation (CAPM) integrates multi-scale visual context with linguistic features using learnable prompts, the language-guided feature aggregation (LGFA) captures cross-scale dependencies to enhance feature aggregation, and the mutual-interaction decoder (MID) ensures precise vision-language alignment through cascaded bidirectional cross-attention, collectively enabling accurate segmentation mask predictions.
@@
    To advance RRSIS research, we construct RISBench, the largest and most diverse benchmark dataset to date, consisting of 52,472 high-quality image-language-label triplets. RISBench features detailed referring expressions, diverse remote sensing scenarios, and pixel-level masks generated through a robust semi-automatic process, providing a comprehensive resource for model training and evaluation.
@@
The remainder of this paper is structured as follows: Section 2 reviews related works on RRSIS. Section 3 details the construction process of our proposed RISBench dataset and provides an analysis of its key characteristics. In Section 4, we describe the proposed methodology in detail. Section 5 presents a comprehensive set of experiments and in-depth analyses. Finally, Section 7 concludes the paper and offers insights into potential future research directions.
2 Related Work
2.1 Referring Image Segmentation
@@
Hu et al.[15] proposed a CNN-LSTM framework for referring image segmentation, effectively extracting visual and linguistic features for precise segmentation. A recurrent refinement network (RRN)[26] was proposed to capture multi-scale semantics in image representations. The RRN iteratively optimized the initial mask using a recursive optimization module to achieve a high-quality pixel segmentation mask.
However, these methods often neglect the modality gap and fail to fully capture the interaction between language and images. Recent works have introduced attention mechanisms to address these limitations. A cross-modal self-attention (CMSA) module by Ye et al.[52] was proposed to effectively captures long-range dependencies between language and visual features. A cascade-grouped attention network (CGAN)[35] is designed, consisting of cascade-grouped attention (CGA) and instance-level attention loss (ILA). By performing hierarchical reasoning on images and effectively distinguishing different instances, CGAN enhances the correlation between text and images. Besides, Hu et al.[17] introduced a bidirectional relationship inferring network (BRINet) to model cross-modal information dependencies. BRINet utilized a visual-guided language attention module to filter out irrelevant regions and enhance semantic matching between target objects and expressions.
2.2 Visual Grounding for Remote Sensing Data
@@
2.3 Referring Remote Sensing Image Segmentation
@@
3 Dataset Construction and Analysis
In this section, we will introduce the construction procedure and statistical analysis of our proposed RISBench dataset in Section 3.1 and Section 3.2.
3.1 RISBench Dataset Construction
  • 1.
  • 2.
    Step 2. To ensure the accuracy and reliability of the generated mask annotations, we employed the PA-SAM method [48] to optimize the SAM model. Specifically, we trained a specialized adapter using a subset of remote sensing images with manually corrected masks, enabling SAM to better handle the unique characteristics of remote sensing imagery, such as complex geospatial relationships and varying object scales. The fine-tuning process involved using bounding box prompts as input, optimizing the mask decoder features with a combination of Dice loss and cross-entropy loss, and iteratively refining the model through multiple rounds of training and validation. After generating the initial masks, we conducted a meticulous human verification process, where expert annotators manually reviewed and corrected the masks. This process included initial inspections, detailed corrections using specialized tools, and a consensus process for uncertain cases, ensuring high-quality and objective annotations.
    Table 1: Comparative analysis of RISBench and prior RRSIS datasets.
    Dataset Triplet Count Image Size Spatial Resolution Expression Attributes
    RefSegRS[55] 4,420 512×512 0.13 m 3
    RRSIS-D[31] 17,402 800×800 0.5 m $\sim$ 30 m 7
    RISBench 52,472 512×512 0.1 m $\sim$ 30 m 8
  • 3.
While potential biases may exist in the dataset due to factors such as annotator subjectivity, dataset characteristics, and linguistic variations, we have taken several steps to minimize their impact. During the annotation process, we employed a consensus-based verification approach, where multiple expert annotators independently reviewed and corrected the masks, ensuring objectivity and accuracy. Additionally, the fine-tuning of the SAM model and the use of diverse remote sensing images helped reduce biases related to object scales and spatial relationships. These measures ensure that the dataset is robust and representative, enhancing the model’s generalization ability to other datasets and real-world scenarios.
@@ 3.1 RISBench Dataset Construction
3.2 Dataset Statistics
After meticulously filtering out uninformative image-language label triplets, we curated the RISBench dataset, comprising 52,472 high-quality image-language label triplets. This dataset is partitioned into a training set with 26,300 triplets, a validation set with 10,013 triplets and a test set with 16,158 triplets, ensuring robust model development and evaluation. Each image in RISBench is uniformly sized at 512$\times$512 pixels, maintaining consistency across the dataset. The spatial resolution of the images spans from 0.1m to 30m, encompassing a diverse range of scales and details. The semantic labels are categorized into 26 distinct classes, each annotated with 8 attributes, thereby facilitating a comprehensive and nuanced semantic segmentation analysis. As shown in Table. 1, compared to the previous RRSIS datasets, our dataset demonstrates significant improvements in both quantity and diversity.
@@ 4 Methodology
4.1 Context-Aware Prompt Modulation
@@
4.2 Language-Guided Feature Aggregation
@@
Subsequently, we introduce an attention deficit compensation mechanism that identifies regions where attention diverges across different scales and explicitly enhances these areas by integrating multi-scale visual features guided by linguistic cues. Specifically, we calculate the attention deficit map $\mathcal{M}\in\mathbb{R}^{H_{4}\times W_{4}}$ between cross-scale attentions and select the top $K$ regions with the highest correlation differences:
@@
4.3 Mutual-Interaction Decoder
@@
Integrating the complementary information between cross-modal features is a fundamental challenge in RRSIS task. Cross-modal features often encompass inconsistent information, and without considering the intermediate interactions and thorough alignment between different modalities, it is impossible to ensure the discriminative power of the learned representations. To address the aforementioned challenges, we have meticulously designed the mutual-interaction decoder (MID), as illustrated in Fig. 8, aiming to achieve more comprehensive vision-language alignment and precise pixel prediction.
The MID utilizes visual features $\left\{V_{il}\right\}_{i=1}^{4}$ and linguistic features $L_{v}$ as inputs to predict the mask of the referred object. Prior to executing vision-language alignment, a series of operations are performed to harmonize the dimensions of the visual and linguistic features.
@@
Unlike existing methods, which typically rely on unidirectional or static interactions that may lead to attention drift or misalignment between modalities, our proposed MID introduces a bidirectional refinement process. This process enables continuous and iterative alignment between visual and linguistic modalities throughout the decoding stage. Such bidirectional interaction ensures that both modalities are dynamically optimized, resulting in more precise feature integration and robust segmentation performance, particularly in complex remote sensing scenarios characterized by diverse geospatial relationships.
4.4 Training Objective
@@ 5 Experiments
5.1 Dataset and Evaluation Metrics
@@
RISBench includes a total of 52,472 image-language-label triplets. It is divided into three subsets: a training set consisting of 26,300 triplets, a validation set consisting of 10,013 triplets and a test set comprising 16,159 triplets. All images are uniformly formatted to a resolution of 512$\times$512 pixels, with spatial resolutions ranging from 0.1 meters to 30 meters. The dataset’s semantic labels are divided into 26 unique classes, with each class further annotated by 8 attributes.
@@
5.2 Experimental Setup
We employ Swin Transformer [32] and ConvNeXt [33] as the visual backbones in our approach. The Swin Transformer backbone is initialized with classification weights from the Swin-Base model pre-trained on ImageNet22K [5]. The ConvNeXt backbone, specifically the ConvNeXt-Base and ConvNeXt-Large variants, is pretrained on the GeoSense dataset, which comprises approximately 9 million diverse remote sensing images, using our proposed SMLFR self-supervised learning framework[10]. This pretraining on GeoSense enables the ConvNeXt backbone to learn robust and generalizable feature representations tailored for remote sensing image interpretation. For the language backbone, we use the base BERT model with 12 layers and a hidden size of 768, as available in the HuggingFace library [45]. The maximum sequence length for text descriptions is set to 20 tokens.
@@ 5.2 Experimental Setup
Table 2: Comparison with state-of-the-art methods on the RRSIS-D dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@
Table 3: Comparison with state-of-the-art methods on the RefSegRS dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@
Figure 9: Visualization of segmentation results for CroBIM and comparison methods on the RRSIS-D dataset test set, with yellow circles highlighting areas of improved performance and corresponding IoU scores displayed.
Figure 10: Visualization of segmentation results for CroBIM and comparison methods on the RefSegRS dataset test set, with yellow circles highlighting areas of improved performance and corresponding IoU scores displayed.
Table 4: Comparison with state-of-the-art methods on the RISBench dataset. Optimal and sub-optimal performance in each metric are marked by red and blue.
@@
5.3 Comparison with State-of-the-art Methods
5.3.1 RRSIS-D
To evaluate the effectiveness of our proposed method, we conducted experiments on the RRSIS-D dataset. The comparison results are presented in Table. 2. We compared CroBIM with several LSTM-based, CLIP-based, and BERT-based methods from classical to state-of-the-art, i.e., RRN[26], CSMA[52], LSCM[19], CMPC[18], BRINet[17], CMPC+[30], BKINet[7], ETRIS[50], CRIS[41], LGCE[55], LAVT[51], RMSIN[31], CrossVLT[4], RIS-DMMI[16], robust-ref-seg[47], SLViT[37], CARIS[29], MAFN[38]. Among the compared methods, LGCE, RMSIN, and MAFN are the only three approaches specifically designed for the RRSIS task in the remote sensing domain. All other methods included in our comparison are SOTA algorithms from the computer vision domain. Although not specifically tailored for remote sensing, these methods represent the current best practices in referring image segmentation and have been widely recognized within the computer vision community.
Figure 11: Visualization of segmentation results for CroBIM and comparison methods on the RisBench dataset test set, with yellow circles highlighting areas of improved performance and corresponding IoU scores displayed.
5.3.2 RefSegRS
We further conduct experiments on the RefSegRS dataset to validate the superiority of our proposed framework, and the performance on both the validation and test sets is reported in Table. 3. First, in terms of the Pr@0.5 and Pr@0.6 metrics, CroBIM consistently outperforms other state-of-the-art methods on both the validation and test sets. Notably, CroBIM with ConvNeXt-B as the visual encoder achieves the highest accuracy of 75.89% on the test set for Pr@0.5, significantly surpassing the second-best method, RIS-DMMI, which attained 63.89%. This result demonstrates that CroBIM effectively segments the target regions under lower overlap thresholds. Additionally, CroBIM continues to exhibit a marked advantage in the Pr@0.6 metric, achieving an accuracy of 61.42% on the test set, which is substantially higher than RIS-DMMI’s 44.30%. As the overlap threshold increases, CroBIM maintains robust performance. Specifically, at the more stringent Pr@0.9 threshold, CroBIM attains an accuracy of 6.26% on the test set, further confirming its effectiveness in high-precision segmentation tasks.
5.3.3 RisBench
In our constructed RISBench dataset, the proposed CroBIM model also demonstrates significant performance advantages over competing methods, achieving either the best or second-best results across multiple evaluation metrics, as shown in Table. 4. Notably, CroBIM outperforms the current state-of-the-art models on both the validation and test sets across various threshold levels, including Pr@0.5, Pr@0.6, Pr@0.7, Pr@0.8, and Pr@0.9. These results further validate the effectiveness and robustness of our model. Specifically, the CroBIM model, utilizing the ConvNeXt-B visual encoder and BERT text encoder, achieves the highest test set precision at Pr@0.5, reaching 77.55%, outperforming all other methods. Additionally, at Pr@0.6, Pr@0.7, and Pr@0.8, the model attains precisions of 72.83%, 66.38%, and 55.93%, respectively, significantly surpassing existing methods. Particularly, at the higher threshold of Pr@0.9, CroBIM leads other methods with a precision of 34.07%, demonstrating the model’s strong capability to handle more complex scenarios that demand higher precision.